OpenStudy (anonymous):

What is the solution to the equation 3^1–2x= 2? a)x ≈ –0.292 b)x≈ –0.185 c)x ≈ 0.185 d)x ≈ 0.292

OpenStudy (anonymous):

is that \[3^{1-2x} = 2?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\log_{3}3^{1-2x} = \log_{3} 2\] \[1 - 2x = \log_{3} 2\] \[2x = 1 - \log_{3}2 \] \[x = (1 - \log_{3}2) / 2 \]

OpenStudy (anonymous):

\[\approx.184535\]